OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4.

OpenStudy (anonymous):

no one likes these conic section problems, do they?

OpenStudy (anonymous):

not really :p

OpenStudy (anonymous):

first lets see what it looks like

OpenStudy (anonymous):

ok isn't it pointing up?

OpenStudy (anonymous):



OpenStudy (anonymous):

oh never mind that then

OpenStudy (anonymous):

no i don't think so the directrix is a vertical line, and the focus is 8 units to the left can you guess which way it opens now?

OpenStudy (anonymous):

it opens to the left right?

OpenStudy (anonymous):

yes. which means the \(y\) term is squared

OpenStudy (anonymous):

ok then i have to use the parabola equation right?

OpenStudy (anonymous):

right now you need to find two things, first the vertex

OpenStudy (anonymous):

the vertex is half way between the focus \((-4,0)\) and the directrix \(x=4\)

OpenStudy (anonymous):

is it clear what that is?

OpenStudy (anonymous):

so that would be (0,0)

OpenStudy (anonymous):

yes, which makes life much easier, because instead of being \(4p(x-h)=(y-k)^2\) it is just \[4px=y^2\]

OpenStudy (anonymous):

now we need \(p\) which is the distance between the vertex and the focus, or the distance between the vertex and directrix (since it is halfway, they are the same thing)

OpenStudy (anonymous):

so 4?

OpenStudy (anonymous):

since it opens to the left, we use \(-4\)

OpenStudy (anonymous):

well the distance is 4 you are right, but as it opens to the left we want \(-4\) now just plug that in to \[4px=y^2\] and you will have it

OpenStudy (anonymous):

want to check it to see if it is right?

OpenStudy (anonymous):

i plug the -4 into the p right?

OpenStudy (anonymous):

right you should get \[-16x=y^2\] and here is the check that it is right scroll down and you will see the vertex and the directrix http://www.wolframalpha.com/input/?i=parabola+-16x%3Dy^2

OpenStudy (anonymous):

thank you but i noticed that this was not an option for my answer

OpenStudy (anonymous):

should be some form of that for sure

OpenStudy (anonymous):

like \[16x=y^2\] or \[x=-\frac{1}{16}y^2\] or something

OpenStudy (anonymous):

the options that i have are y^2 = -8x 16y = x^2 y = -(1/16)x^2 x = -(1/16)y^2

OpenStudy (anonymous):

ah yes i just saw your other answer

OpenStudy (anonymous):

last one is the same as what we wrote

OpenStudy (anonymous):

i want to thank you so much for helping me

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i am going another one now if you want to see it this one is a hyperbola

OpenStudy (anonymous):

ah yes that would be nice